Reports 1-1 of 1 Clear search Modify search
AdV-DET (Commissioning)
gouaty - 21:25 Wednesday 18 February 2026 (68715) Print this report
Comment to OMC mode matching tuning (68707)

Looking at the data after the ITF was relocked in Low Noise 3 aligned (attached figure), we can see that there is an offset between the B5 QD2 signal (in particular in H) and the SDB1 bench aligned position with the dither control. I already added an horizontal offset of -170 um in the SDB1_Quadrants configuration file, but the configuration needs to be reloaded to apply it. To be done when the interferometer will be DOWN.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×